### v2

# DefendCLI: {Command-Line} Driven Attack Provenance Examination

Abstract
Endpoint Detection and Response (EDR) solutions embrace the method of attack provenance graph to discover unknown threats through system event correlation. However, this method still faces some unsolved problems in the fields of interoperability, reliability, flexibility, and practicability to deliver actionable results. Our research highlights the limitations of current solutions in detecting obfuscation, correlating attacks, identifying low-frequency events, and ensuring robust context awareness in relation to command-line activities.
To address these challenges, we introduce DefendCLI, an innovative system leveraging provenance graphs that, for the first time, delves into command-line-level detection. By offering finer detection granularity, it addresses a gap in modern EDR systems that has been overlooked in previous research. Our solution improves the precision of the information representation by evaluating differentiation across three levels: unusual system process calls, suspicious command-line executions, and infrequent external network connections. This multi-level approach enables EDR systems to be more reliable in complex and dynamic environments.
Our evaluation demonstrates that DefendCLI improves precision by approximately $1.6\times$ compared to the state-of-the-art methods on the DARPA Engagement Series attack datasets. Extensive real-time industrial testing across various attack scenarios further validates its practical effectiveness. The results indicate that DefendCLI not only detects previously unknown attack instances, which are missed by other modern commercial solutions, but also achieves a $2.3\times$ improvement in precision over the state-of-the-art research work.
I Introduction
Traditional security measures, like anti-virus software and firewalls, frequently fall short in detecting contemporary threats that employ advanced evasion strategies, such as memory injection and fileless attacks. These limitations highlight the need for advanced Endpoint Detection and Response (EDR) solutions to detect complex and stealthy attacks, providing better insights into behavior and operations to improve security incident detection. An essential improvement of modern EDR systems is the attack provenance graph, which visually represents the relationships between different system events to discover potential advanced threats.
While there is growing interest in provenance-based EDR systems, their adoption in real-world scenarios remains limited, and comprehensive evaluation is still lacking. To address this gap, we deployed a provenance-based EDR [33] to automatically analyze system audit logs from the Microsoft Cyber Defense Operations Center (CDOC) and passed the results to security analysts for verification. During this process, analysts encountered challenges in efficiently utilizing the provenance-based results and we summarized four key unresolved research questions (RQs) that require further investigation.
Figure 1: Motivation Feedback: A provenance-based detector [33] deployed in an enterprise environment. Only 0.07% of the original alerts were truly actionable, providing accurate attack-related malicious activities, such as command-line executions and network connections, for a valid verification.
1
RQ1-Interoperability: How can we reduce the complexity of attack provenance graphs to provide more actionable results? Provenance graphs are often overly complex and cluttered with invalid information, making interpretation difficult and misaligned with analysts’ priorities.
Malicious activities often blend seamlessly into legitimate system operations, causing provenance-based EDRs to generate an overwhelming number of anomalies that lack actionable insights, as illustrated in Fig 1 and Fig 2. While these results may not necessarily be false alarms, they fail to provide definitive evidence of an intrusion, making manual investigation even more challenging and hindering efforts to trace the root cause and focus on suspicious activities.
These challenges highlight the need for a more refined approach to provenance graph construction, which can minimize complexity, enhance detection accuracy, and deliver more relevant, actionable insights for security analysts.
Figure 2: Omission: Even if a process chain is malicious, but it cannot be analyzed if we do not known what command-line are executed, of which effective threat verification is affected.
2
RQ2-Reliability: How can we provide a more effective approach for anomaly detection on different operating system platforms? Existing anomaly detection methods [9, 34, 8] often adopt a one-size-fits-all approach, failing to account for OS-specific characteristics. While this generalization increases versatility, it overlooks critical real-world complexities.
Enterprise security policies often segment networks into distinct zones, such as the demilitarized zone (DMZ) and internal office networks. The DMZ primarily hosts externally facing services, like DNS and web servers that typically run Linux, whereas internal office networks support end-user devices, predominantly Windows-based. Each segment presents unique behavioral patterns, attack surfaces, and vulnerabilities, making OS-specific anomaly detection essential for effective threat identification.
A more tailored approach, one that considers the distinct operational norms and security risks of each OS, can significantly improve the accuracy and efficiency of EDR solutions. By refining detection mechanisms to align with platform-specific behaviors, EDR systems can reduce false positives, enhance threat visibility, and provide more actionable results.
3
RQ3-Flexibility: How can we tailor AI-based detection mechanisms to minimize false positives and improve adaptability in dynamic environments? The adoption of AI in security detection and response is accelerating due to its superior capability in identifying zero-day threats compared to traditional signature-based systems.
However, AI-driven solutions introduce significant challenges, particularly an elevated false positive rate in dynamic environments where factors such as new device deployments, software updates, and network reconfigurations frequently trigger erroneous alerts.
Moreover, the inherent “black-box” nature of AI complicates security investigations, as opaque decision-making processes undermine interpretability, trust, and operational transparency.
To address these limitations, AI models should incorporate adaptive learning mechanisms that accommodate environmental variability while maintaining a stringent false positive control. Enhancing explainability through interpretable AI techniques and leveraging contextual awareness in anomaly detection can further improve precision, thereby fostering greater trust and operational efficacy in security operations.
4
RQ4-Practicability: How can we evolve beyond binary decision-making frameworks in EDR systems to enable more granular and nuanced threat analysis? Traditional security frameworks that classify alerts as strictly malicious or benign are increasingly inadequate, as modern cyber threats often blur the lines between legitimate and malicious activity. Advanced attack techniques exploit these gray areas, rendering binary classifications ineffective in identifying sophisticated threats.
To address this limitation, AI-driven security solutions are shifting toward threat-scoring models that assess incidents based on severity, providing a more nuanced analysis. However, many of these systems remain rule-based, necessitating frequent updates to adapt to evolving attack patterns. This continuous maintenance imposes a significant operational burden on security vendors, limiting scalability and responsiveness to emerging threats.
To enhance efficiency, a more automated, adaptive and actionable security orchestration framework is required - one that minimizes manual intervention while dynamically refining threat assessments. By integrating machine learning-driven anomaly detection, contextual analysis, and automated incident response, EDR systems can improve detection precision, reduce operational overhead, and offer a more intelligent, scalable approach to threat management.
Our Contributions. To address the aforementioned RQs, DefendCLI presents a practical solution that, for the first time, delves into the command-line-level detection within the attack provenance graph, improving accuracy while avoiding unrelated information.
Command-line interactions encompass both interactive shell operations and OS API executions, providing direct visibility into system-level behaviors. Despite attackers’ attempts at command-line obfuscation, sophisticated multi-stage attacks require numerous system function executions, inevitably producing traceable command-line artifacts that cannot be fully concealed from detection systems. Despite the widespread adoption of command-line monitoring in rule-based EDR systems, its formal representation in attack provenance graphs and systematic integration with AI-driven anomaly detection frameworks remains underdeveloped.
Our solution employs an explainable AI system to match the known malicious command-line with well-defined rule sets and detect unknown command-line activities with the SimHash-Based [27] anomaly detection mechanism. Furthermore, the overall solution focuses on multi-level attack provenance examination, including unusual system process calls, suspicious command-line executions, and infrequent external network connections, which can resist defense evasion to some extent.
To provide a more automated security orchestration scheme, we integrate Large Language Models (LLMs) into our threat recommendation system to generate rapid preliminary assessments, enabling efficient triage before comprehensive threat analysis. Extensive experiments and evaluations show that DEFENDCLI effectively manages real-world cyber threats by identifying previously undetected attack instances.
Our contributions are summarized as follows:
  • Interoperability: We introduce a novel attack provenance graph structure specifically tailored for command-line-driven, multi-level differentiation detection. This refined structure reduces the complexity associated with traditional provenance graphs, resulting in more precise information representation and improved actionable insights for security analysts.
  • Reliability: Our approach incorporates an explainable anomaly detection framework, integrating defense evasion signatures and known attack patterns from both Linux and Windows platforms into an AI-driven architecture. The system also employs graph algorithms for multi-level provenance analysis alongside SimHash-based differentiation of unknown command-line executions, significantly reducing false positives by moving away from pre-training unreliable normal behavior profiles.
  • Flexibility: The framework supports dynamic adaptation through configurable hyperparameters, optimizing performance for various operational environments. By leveraging a retrieval-augmented generation (RAG) approach with LLMs, we generate prioritized incident assessments with detailed explanations, named snapshot alarms, thus streamlining alert triage and mitigating analyst alert fatigue.
  • Practicability: We evaluate the performance of our system against state-of-the-art research work and modern commercial EDR tools on academic benchmark datasets and industrial real-time detection. The result shows the practical applicability and superior effectiveness of our solution for modern offensive and defensive confrontation.
II Related Work
This section provides a comparative analysis of related work across four key dimensions critical for command-line activity detection: obfuscation detection (OD), attack correlation (AC), low-frequency detection (LFD), and context awareness (CA) [23]. Table I presents a feature analysis of DefendCLI relative to prior methods based on their algorithmic designs (AD).
1
OD: Challenges in Obfuscation Detection [6]. Most existing approaches primarily leverage structural anomalies, a strategy that may not fully capture the semantics of obfuscated commands (e.g., polymorphic commands or encoded payloads). Holmes [22] provides effective detection via TTP-based rules. Methods like Unicorn [7], DepComm [35], ProGrapher [36], Nodlink [17], and DISTDET [5] focus largely on structural changes or statistical deviations. While effective for general anomalies, these systems face challenges in parsing sophisticated obfuscation techniques without semantic analysis.
2
AC: Granularity in Attack Correlation [37]. Effective correlation is essential for reconstructing attack chains. Nodlink [17] demonstrates strong performance by employing Steiner trees for node-level correlation, while Holmes [22] utilizes TTP rules to provide a cohesive view of adversarial behavior. Unicorn, DepComm, and ProGrapher rely on dependency tracking, which captures structural relationships. DISTDET [5] emphasizes efficient log reduction and deduplication, which prioritizes scalability but may limit the resolution of complex relationships in multi-stage APTs.
3
LFD: Sensitivity to Low-Frequency Events [24]. Detecting rare, non-repetitive patterns is vital for identifying novel or targeted attacks. Many current methods (Unicorn, DepComm, ProGrapher, Nodlink, DISTDET) utilize historical frequency profiles, which are effective for repetitive behaviors but may be less sensitive to rare events lacking precedence. Holmes [22] offers detection through predefined rules, whereas anomaly-based baselines in other systems must balance sensitivity with false-positive rates when encountering stealthy, low-frequency threats.
4
CA: Context Awareness Integration [18]. Distinguishing benign from malicious commands often depends on execution context (e.g., user privileges, parent processes). DepComm [35] incorporates system attributes as dependencies to enhance context. Holmes [22] uses rule-based context for known attacks. Approaches such as Unicorn, ProGrapher, Nodlink, and DISTDET primarily focus on graph structure. Integrating robust environmental variables remains a key area for development to accurately differentiate between legitimate operations and malicious mimicry.
Figure 3: Attack-Clause Sketch and Attack-Evidence Awareness: The attack-clause sketch module includes (1) provenance construction for building behavior graphs and (2) command-line evaluation to refine weights using known attacks. The attack-evidence awareness module handles (3) anomaly detection for unknown threats and (4) reporting via RAG-LLM.
TABLE I: Feature Analysis of DEFENDCLI Relative to Prior Work. : Comprehensive, : Substantial, : Partial, : Limited.
Method OD AC LFD CA AD
Holmes [22] ATT&CK Framework
Unicorn [7] Path Traversal
DepComm [35] Graph Partition
Prographer [36] Graph Partition
Nodlink [17] Steiner Tree Partition
DISTDET [5] Graph Partition
DefendCLI Multi-Level Deviation
III DefendCLI
The architecture of DefendCLI, illustrated in Figure 3, adopts a highly modular design to ensure extensibility and clarity in addressing the challenges of command-line attack detection. The system is fundamentally composed of two sequential computational phases: Attack-Clause Sketch, dedicated to the construction, refinement, and structural quantification of the provenance graph; and Attack-Evidence Awareness, responsible for semantic analysis, anomaly detection, and explainable threat reporting.
III-A Attack-Clause Sketch: Graph Construction and Quantification
This module focuses on transforming raw system telemetry into a structured, quantifiable representation suitable for causal analysis. It involves graph refinement, probabilistic node scoring, and interphase community association.
Figure 4: Comparison of (a) Information-Flow Focused Structure (heterogeneous nodes) and (b) Refined Command-Line Focused Structure (isomorphic process nodes with attributes). The refined structure reduces graph noise by encapsulating command and network data within process nodes.
Figure 5: Attack-chain causal inference: Node importance is calculated using PageRank (for rarity) and Shortest-Path Betweenness Centrality (for structural bridges). These scores are aggregated to determine initial edge weights.
Figure 6: Interphase attack association: The Leiden community detection algorithm groups related process nodes, overcoming fragmentation issues to correlate multi-stage attacks.
III-A1 Refined Provenance Structure and Reduction
Traditional approaches often construct heterogeneous graphs where processes, files, sockets, and command strings are distinct nodes (Figure 4(a)). This results in high-dimensional, noisy graphs where semantic context is diluted.
To address this, we define a Refined Command-Line Focused Structure (Figure 4(b)). Let the system execution be represented as a directed graph $G=(V,E,\mathcal{A})$. Here, $V$ represents the set of uniform system process nodes. $E\subseteq V\times V$ denotes direct causal dependencies (e.g., process forking). Crucially, command-line arguments and network contexts are not separate nodes but are encapsulated as an attribute set $\mathcal{A}_{v}=\{\text{cmd}_{v},\text{net}_{v},\dots\}$ resident within each process node $v\in V$. This isomorphic structure preserves semantic richness while significantly reducing graph complexity for downstream embedding and traversal algorithms.
To mitigate the dependency explosion problem [10], we apply a two-stage graph reduction function $\Phi:G\to G^{\prime}$.
First, Semantic Pruning removes nodes lacking command-line relevance. We define a subgraph $G_{sub}\subset G$ as irrelevant if for all nodes $v\in G_{sub}$, the command attribute $\text{cmd}_{v}=\emptyset$. These subgraphs are pruned as they typically represent benign background noise.
Second, Acyclic Transformation ensures valid causal tracing by eliminating cyclic dependencies. $G^{\prime}$ is transformed into a Directed Acyclic Graph (DAG) via depth-first search based cycle removal, ensuring that any traversal represents a valid temporal flow of events.
III-A2 Causal Inference via Hybrid Node Scoring
Identifying critical nodes within massive provenance graphs requires quantifying their structural importance. We employ a hybrid scoring mechanism combining probabilistic ranking and centrality measures to determine initial node weights, as depicted in Figure 5.
Quantifying Rareness (Inverted PageRank). We adapt the PageRank algorithm [2] to identify anomalous nodes based on topological scarcity. Let $PR(v)$ denote the PageRank score of node $v$. The iterative computation is defined as:
$PR(v)_{t+1}=(1-d)+d\sum_{u\in\text{In}(v)}\frac{PR(u)_{t}}{L(u)}$ (1)
where $d\approx 0.85$ is the damping factor, $\text{In}(v)$ are incoming predecessors, and $L(u)$ is the out-degree of $u$. Unlike standard applications where high PR indicates prominence, in attack provenance, nodes with low in-degree (lower PR) often signify rare, potentially malicious execution paths. We normalize these scores to emphasize rarity.
Identifying Root Causes (Betweenness Centrality). To find pivotal ”bridge” nodes that connect disparate attack phases (e.g., initial access leading to privilege escalation), we utilize Shortest-Path Betweenness Centrality ($CB$) [3]. For a node $v$, $CB(v)$ quantifies the fraction of shortest paths between all node pairs $(s,t)$ that pass through $v$:
$CB(v)=\sum_{s\neq v\neq t}\frac{\sigma_{st}(v)}{\sigma_{st}}$ (2)
where $\sigma_{st}$ is the total number of shortest paths from $s$ to $t$, and $\sigma_{st}(v)$ is the count of those paths passing through $v$. High $CB$ scores highlight structurally critical bottlenecks.
Edge Weight Initialization. To facilitate path-based analysis, these node-level metrics must be projected onto the edges. For any directed edge $e_{mn}\in E^{\prime}$ connecting parent $m$ to child $n$, we define the initial Base Edge Weight ($W_{mn}^{base}$) as the normalized aggregate of the structural significance of both connected nodes:
$W_{mn}^{base}=\text{Norm}\left(PR(m)+PR(n)+CB(m)+CB(n)\right)$ (3)
where $\text{Norm}(\cdot)$ scales the resulting magnitude to the range $[0,1]$. This weight represents the inherent structural interest of the transition from $m$ to $n$.
III-A3 Interphase Attack Association via Community Detection
Attacks often manifest as distinct phases separated by time, making causal linkage difficult. To correlate these multi-stage behaviors, we partition the reduced graph $G^{\prime}$ into semantically coherent clusters.
We utilize the Leiden algorithm [16], an improvement over Louvain that guarantees connected communities and optimizes modularity. As illustrated in Figure 6, Leiden partitions the graph into a set of disjoint communities $\mathcal{C}=\{C_{1},C_{2},\dots,C_{k}\}$. Each detected community $C_{k}$ is assigned a Community Score ($S_{C_{k}}$) based on the aggregate risk density within that cluster. This score serves as a coherence multiplier in subsequent analysis, reinforcing links between activities occurring within the same attack phase context.
III-B Attack-Evidence Awareness: Detection and Reporting
The second phase leverages the weighted provenance graph to detect threats. This involves differentiating command-line semantics against known signatures and employing ensemble learning for unknown anomaly detection, as outlined in Figure 7.
Figure 7: Explainable anomaly detection mechanism: The process involves retrieving ranked InfoPaths, analyzing command-line differences via SimHash, and using a multi-classifier ”Police-Witness” voting policy for final verdict.
III-B1 Differentiation of Known Threats (Rule-Based Boosting)
To integrate expert knowledge, we evaluate process command-lines against curated threat intelligence, specifically Sigma rules [29] and defense evasion patterns [12]. Let $R(e_{mn})$ be a boolean indicator function that returns 1 if the command-line context $\text{cmd}_{n}$ associated with edge $e_{mn}$ matches a known threat signature, and 0 otherwise. We introduce a Risk Coefficient $\alpha>1$ corresponding to the severity of the matched rule. The base edge weight is updated to a Refined Edge Weight ($W_{mn}^{refined}$) to reflect this deterministic knowledge:
$W_{mn}^{refined}=\begin{cases}W_{mn}^{base}\times\alpha&\text{if }R(e_{mn})=1\\ W_{mn}^{base}&\text{otherwise}\end{cases}$ (4)
This operation significantly amplifies the weights of paths containing known malicious indicators, ensuring their prioritization during retrieval.
III-B2 Differentiation of Unknown Anomalies (SimHash Ensemble)
Detecting novel attacks requires identifying subtle deviations in command-line semantics, often disguised through obfuscation. We propose an approach combining semantic embedding with locality-sensitive hashing.
Semantic Hashing via SimHash. Standard embeddings (e.g., Word2Vec) result in high-dimensional vectors where similarity search is computationally expensive. To efficiently detect obfuscated variations, we apply SimHash [27]. Let $\mathbf{v}_{c}\in\mathbb{R}^{d}$ be the dense vector representation of a command string $c$. The SimHash function $H:\mathbb{R}^{d}\to\{0,1\}^{f}$ maps this vector to an $f$-bit binary signature $h_{c}$. A crucial property of SimHash is that the Hamming distance between two signatures approximates the cosine similarity of their original vectors:
$\text{Hamming}(h_{c_{i}},h_{c_{j}})\approx\theta(\mathbf{v}_{c_{i}},\mathbf{v}_{c_{j}})$ (5)
This allows us to efficiently quantify semantic deviation, as shown in Figure 8, where obfuscated commands yield small but detectable Hamming distances from benign counterparts.
Figure 8: Example of SimHashed command-line executions. SimHash maps semantically similar commands to nearby binary signatures, allowing detection of obfuscated variations (red box) via Hamming distance.
Ensemble ”Police-Witness” Voting. To avoid algorithmic bias inherent in single-model anomaly detection, we implement an ensemble voting strategy [25]. We employ a diverse set of unsupervised classifiers $\mathcal{M}=\{\text{iForest},\text{LOF},\text{OC-SVM},\text{EllipticEnvelope},\dots\}$. Each classifier $M_{j}\in\mathcal{M}$ evaluates the SimHashed InfoPaths independently. The final classification of a path $P$ is determined by a majority vote:
$V(P)=\mathbb{I}\left(\sum_{M_{j}\in\mathcal{M}}\mathbb{I}(M_{j}(P)=\text{anomalous})>\frac{|\mathcal{M}|}{2}\right)$ (6)
Paths confirmed as anomalous by the majority are flagged for analyst review.
III-B3 InfoPath Retrieval and GPT-Powered Reporting
The final step is to extract the most salient attack chains, termed InfoPaths, from the graph. An InfoPath is defined as a directed sequence of causally linked process nodes $(v_{1},v_{2},\dots,v_{k})$ representing a potential attack progression.
To retrieve the most critical paths, we adapt Dijkstra’s shortest-path algorithm [13]. Since standard Dijkstra minimizes cost, and our refined weights $W_{mn}^{refined}$ represent *evidence* (where higher is more significant), we define the traversal cost function as the inverse of the evidence, further weighted by the community score to favor cohesive attack phases:
$Cost(e_{mn})=\frac{1}{W_{mn}^{refined}\times S_{C_{community(n)}}}$ (7)
By minimizing this cost, the algorithm effectively maximizes the cumulative evidence along the path, prioritizing InfoPaths with high structural rarity, centrality, known threat indicators, and community coherence.
To mitigate alert fatigue, the Top-$k$ retrieved InfoPaths are fed into a Retrieval-Augmented Generation (RAG) framework [26] powered by Llama-2. The model augments the raw path data with context from external threat intelligence reports, generating a concise, prioritized snapshot alarm that explains the attack intent and recommended actions.
IV Evaluation
The evaluation is divided into overall evaluation and ablation evaluation. The overall evaluation compares our approach with state-of-the-art methods using precision, recall, and $F_{1}$-score on selected academic and industrial attack datasets. The ablation evaluation conducts a detailed functional analysis to assess the effectiveness of each component in our solution, providing deeper insights into its impact and performance on discussed RQs and limitations.
IV-A Attack Datasets
IV-A1 DARPA Engagement Series Datasets (Linux and Unix-Like Servers)
The DARPA Transparent Computing Engagement Series (TCES) [4] are designed to support cybersecurity research by providing real-world cyber attack traces, system logs, and network activity. Due to most of related work are evaluated on the Engagement-3 (E3) dataset, to make a fair comparison, we also use the E3 dataset as the benchmark. The E3 dataset captures multi-user cyber attack scenarios with a mix of legitimate and malicious activities, including privilege escalation, lateral movement, and data exfiltration, making it ideal for anomaly detection and forensic analysis. It includes host-level logs, file system events, system calls, and process activities, making them valuable for developing machine learning models, behavioral analytics, and proactive cyber defense strategies. Based on the attack scenarios, the E3 dataset is divided into E3-Trace, E3-THEIA and E3-CADETS.
TABLE II: GitHub Availability, Reproducibility and Real-Time Detection of Related Work
Method A RP RT C
Holmes [22] S&P
Unicorn [7] NDSS
DepComm [35] S&P
Prographer [36] Usenix Security
Nodlink [17] NDSS
DISTDET [5] Usenix Security
TABLE III: Licensed Commercial EDR Solutions
Tool AV WS CTI RT
Microsoft Defender [20]
Symantec EDR [32]
Karspersky EDR [14, 15]
TABLE IV: Node-Level Results on DARPA E3 Datasets
Method E3-Trace E3-THEIA E3-CADETS
Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score
Holmes [22] $1.35\times 10^{-3}$ 0.74 $2\times 10^{-3}$ $3.61\times 10^{-3}$ 0.98 0.01 $2.84\times 10^{-3}$ 0.95 0.01
Unicorn [7] $3.20\times 10^{-5}$ N/A N/A $1.86\times 10^{-4}$ N/A N/A $1.25\times 10^{-4}$ N/A N/A
DepComm [35] $2.3\times 10^{-3}$ 1 $4.6\times 10^{-3}$ $1.2\times 10^{-2}$ 1 0.02 $1.2\times 10^{-2}$ 1 0.02
Nodlink [17] 0.25 0.98 0.40 0.23 1 0.37 0.14 1 0.25
DefendCLI 0.41 1 0.58 0.38 1 0.55 0.2 1 0.33
TABLE V: Attack Scenario Checklist (E3)
Dataset Attack Scenario TTPs
TRACE Firefox Exploit via Ads Drive-by Execution
Malicious Extension Drops Payload DLL Injection
Phishing Email Installs Backdoor Credential Theft
THEIA Firefox Exploit Injects Script Remote Execution
Fake Extension Steals Passwords Keylogging Attack
CADETS Web Server Exploited Remotely Process Injection
Fake Update Installs Malware Persistence Attack
Lateral Movement via Credentials Privilege Escalation
(a) Attack-Related Command-Line (E3-TRACE)
(b) Attack-Related Command-Line (E3-THEIA)
(c) Attack-Related Command-Line (E3-CADETS)
(d) InfoPath-Level Result (E3-TRACE)
(e) InfoPath-Level Result (E3-THEIA)
(f) InfoPath-Level Result (E3-CADETS)
Figure 9: Comparison of attack-related command-line and InfoPath-Level results on E3 datasets.
IV-A2 Industrial Real-Time Detection (Microsoft Windows End-User and Servers)
The Industrial Real-Time Detection dataset [11] is designed to support cybersecurity research by providing real-world, real-time telemetry from virtual machines (VMs) hosted on Microsoft Azure, enabling the evaluation of real-time detection and response capabilities for defense solutions. The dataset captures system logs, network activity, and security events in an industrial environment, simulating active cyber threats and operational challenges as they unfold. By including both benign and malicious activities, it allows researchers to study and enhance real-time threat detection, anomaly identification, and rapid intrusion response in cloud-based infrastructures. With realistic attack scenarios such as unauthorized access, privilege escalation, and lateral movement, this dataset is particularly valuable for advancing machine learning-driven security solutions, continuous monitoring systems, and proactive cyber defense strategies, emphasizing the importance of real-time detection and mitigation in industrial cloud environments.
TABLE VI: Attack-TTPs and Node-Level Results on Industrial Real-Time Detection
Method Attack Scenario #1 (A1) Attack Scenario #2 (A2) Attack Scenario #3 (A3)
Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score Precision Recall $F_{1}$ Score
Microsoft Defender [20] 0.04 1 0.08 0 1 0 0 1 0
Symantec EDR [32] block block block 0.13 1 0.23 0.04 1 0.08
Kasparsky EDR [14] block block block block block blcok 0.09 1 0.10
Nodlink [17] (TTPs) 1 1 1 1 1 1 1 1 1
Nodlink [17] (Node-Level) 0.04 1 0.08 0.19 1 0.32 0.05 1 0.10
DefendCLI (TTPs) 1 1 1 1 1 1 1 1 1
DefendCLI (Node-Level) 0.08 1 0.15 0.38 1 0.55 0.18 1 0.31
TABLE VII: Attack Scenario Checklist (Industrial)
Scenario APT TTPs
1 Silent Lynx [30, 31] PasteBin Multi-Staged Powershell
2 DriftingCloud [1, 19] MITM PHP Chopper Webshell
3 Evilnum [28] Spear Phishing Python RAT
IV-B GitHub Availability, Reproducibility and Real-Time Detection
Table II shows the availability (A), reproducibility (RP), and real-time detection capability (RT) of related work based on big-4 security conferences (C). The research work of DISTDET [5] and Prographer [36] are not publicly available on GitHub or other sources. Therefore, our comparative study focuses on the remaining work. The verification of availability was conducted on 7 February 2025.
IV-C Licensed Commercial EDR Solutions
Table III presents the detection capabilities of commercial EDR solutions across four key security features: anti-virus (AV), web security (WS), cyber threat intelligence (CTI), and real-time (RT) protection. For our evaluation, we obtained commercial licenses for Microsoft Defender [20], Symantec EDR [32], and Kaspersky EDR [14, 15]. Each EDR solution was updated to the latest version, ensuring the most recent threat database was used during testing.
IV-D Attack Provenance Evaluation Criterion
We use a multi-level criterion to evaluate our work, which is introduced as follows:
  • InfoPath-Level: The InfoPath-level evaluation criterion was introduced by DefendCLI, which detects anomalous attack chains by analyzing multiple paths between source and destination nodes. InfoPaths represent the interactions between system behaviors and serve as a core metric in both real-time industrial detection scenarios and our ablation studies.
  • Node-Level The node-level criterion, first introduced in Nodlink [17], addresses the limitations of prior approaches that evaluated entire attack provenance graphs often achieving perfect precision while inadvertently including many non-attack-related events. Instead, this criterion focuses on identifying and quantifying attack-relevant nodes within InfoPaths, offering a more granular and accurate reflection of actual attack activity.
  • TTPs-Level The TTPs-level criterion is used to compare detection capabilities against commercial EDR tools. Since many EDRs terminate execution upon detecting an attack, limiting further behavior observation, we assess the number of detected activities associated with adversarial tactics, techniques, and procedures (TTPs). This allows for a direct performance comparison between our solution and commercial detection tools.
IV-E Overall Evaluation
IV-E1 Detection Results on DARPA E3 Benchmark
Table V presents the threat model of the three datasets in the E3 benchmark, while Table IV compares node-level detection results with state-of-the-art research. Our analysis shows that while all existing methods can detect E3 attack scenarios, their node-level results vary significantly. This discrepancy highlights the presence of substantial non-attack-related information in their alarms, leading to lower detection precision and reduced $F_{1}$ scores.
Our solution DefendCLI, improves detection by focusing on command-line-level analysis, achieving a precision improvement of 1.6 $\times$ over the best existing method Nodlink [17]. Figure 9 presents the results at the InfoPath-level and the command-line activities related to the detected attack identified by DefendCLI. This improvement demonstrates that DefendCLI provides more concise results, detecting E3 attack scenarios with fewer nodes while delivering more actionable insights for threat incident verification, significantly enhancing detection efficiency.
(a) Attack-Related Command-Line (Real-Time)
(b) Real-Time InfoPath Alarm (Attack Scenario #1)
(c) Real-Time InfoPath Alarm (Attack Scenario #2)
(d) Real-Time InfoPath Alarm (Attack Scenario #3)
Figure 10: Comparison of attack-related command-line and InfoPath alarms in real-time detection.
IV-E2 Detection Results in the Industrial Real-Time Detection
Table VII presents the threat model for industrial real-time detection, incorporating modern APT threats relevant to 2025. These advanced threats are designed to evade modern defense strategies, enabling a more comprehensive evaluation of our solution’s detection capabilities in complex industrial environments.
[…]
Figure 10 illustrates the attack-related command-line activities and the number of real-time InfoPath alarms identified by DefendCLI. The industrial environment presents greater complexity than the E3 datasets due to a significant number of active users generating normal activities during evaluation. Despite this, our results show that DefendCLI maintains a sound alarm rate relative to the volume of generated events, demonstrating its practicality for real-world deployment.
(a) Attack-Chain Causal Inference (E3-TRACE)
(b) Attack-Chain Causal Inference (E3-THEIA)
(c) Attack-Chain Causal Inference (E3-CADETS)
(d) Attack-Chain Causal Inference (Attack Scenario #1)
(e) Attack-Chain Causal Inference (Attack Scenario #2)
(f) Attack-Chain Causal Inference (Attack Scenario #3)
Figure 11: Comparison of attack-chain causal inference based on the number of weighted edge count versus the number of attack-related weighted edge count.
IV-F Ablation Evaluation
IV-F1 Evaluating the Interoperability of Attack Provenance Examination Through Attack-Clause Sketch and Evidence Awareness
Our solution incorporates a variety of sophisticated algorithms to enhance the detection capabilities of attack provenance analysis. In this section, we present a detailed assessment, breaking down the contribution of each module to demonstrate their individual effectiveness within our framework.
$\blacksquare$ How well does attack-chain causal inference identify unusual elements within the attack provenance graph? The effectiveness of attack-chain causal inference in detecting unusual elements within the attack provenance graph is evaluated based on the comparison between the total number of weighted edges and the subset of edges directly associated with attack activities. Figure 11 illustrates this comparison in different scenarios, highlighting the ability of our method to isolate suspicious elements from normal system interactions.
The results demonstrate that our causal inference model successfully assigns precise and meaningful weights to attack-related edges, ensuring that critical attack components are effectively highlighted while enabling efficient path traversal. This mechanism allows for a clear distinction between malicious activities and benign system interactions, minimizing noise, and improving detection accuracy. The observed disparity between the total number of weighted edges and the subset associated with attacks further validates the model’s ability to prioritize attack-relevant edges, effectively filtering out irrelevant connections and improving the interpretability of the attack provenance graph.
TABLE VIII: The Use and Not Use of Community Detection and SimHash Recommender on Node-Level $F_{1}$ Score
Dataset Community Detection SimHash (Blue)
Y N W2V D2V FT
E3-TRACE 0.58 0.47 0.58 0.48 0.50
E3-THEIA 0.55 0.54 0.31 0.42 0.55
E3-CADETS 0.33 0.30 0.33 0.31 0.32
A1 0.15 0.13 0.14 0.14 0.15
A2 0.55 0.31 0.48 0.30 0.55
A3 0.31 0.31 0.31 0.18 0.26
$\blacksquare$ How successfully does interphase attack association enhance the correlation of multi-staged attacks? Figure 12(a) compares the use of community detection in analyzing the provenance of the attack, where Y/N denotes whether community detection is applied. The blue result highlights the better outcome. We reconstructed the detected InfoPaths as weakly connected attack provenance subgraphs to assess how many reachable elements could be accurately correlated.
For the E3-TRACE and E3-THEIA datasets, we observed that without community detection, our solution results in a higher false alarm rate (FAR) while attempting to identify all attack activities. In contrast, for the E3-CADETS dataset, particularly in industrial attack scenarios #1 and #2, we found that while the absence of community detection may lead to fewer results, some attack activities go undetected. However, for industrial attack scenario #3, we observed no significant difference, as all elements were reachable regardless of the use of community detection.
Table VIII shows the node-level $F_{1}$ score with or without the use of community detection. The experimental results can support our claim that the application of community detection improves correlation accuracy for identifying interphase attacks, leading to more precise and effective threat analysis.
$\blacksquare$ How proficient is the SimHash recommender in distinguishing unknown command-line activities? Figure 12(b) compares the use of the SimHash recommender in analyzing the provenance of the attack, where W2V, D2V, and FT represent the use of Word2Vec, Doc2Vec, and FastText, respectively. The blue result highlights the selection of the SimHash recommender. For the E3-TRACE, E3-CADETS and attack scenario #3, SimHash recommender selected Word2Vec as the method of vectorization because it has the maximum score to distinguish the command-line differentiation. On the other hand, for the E3-THEIA, attack scenario #1 and #2, SimHash recommender selected FastText as the method of vectorization based on the same standard.
Similarly, we observed that for E3-TRACE and attack scenarios #2 and #3, SimHash generated fewer results while still detecting all attack behaviors, effectively reducing the FAR. In contrast, for E3-THEIA, E3-CADETS, and attack scenario #1, SimHash produced a larger set of results, ensuring comprehensive attack detection. However, other methods, despite yielding fewer results, failed to identify several malicious behaviors.
Table VIII presents the node-level $F_{1}$ score with and without the use of SimHash recommender. The results demonstrate that incorporating the SimHash recommender enhances context awareness by effectively differentiating command-line activities, leading to more precise threat analysis.
(a) Community Detection
(b) SimHash Recommender
Figure 12: Comparison of the use with or without community detection and SimHash recommender based on the count of InfoPaths detected.
(a) Diverse OS Platforms
(b) Adaptive Interface
Figure 13: Comparison of the $F_{1}$ scores on diverse OS platforms and the use of adaptive interface.
IV-F2 Cross-Platform Detection Reliability: Performance Across Heterogeneous OS Environments
Figure 13(a) shows the node-level $F_{1}$ score across different OS environments. In our evaluation, E3-TRACE and E3-THEIA were conducted on a Linux system, E3-CADETS was based on a FreeBSD system, and attack scenarios #1 to #3 were evaluated on a Windows system. It can be seen that our solution benefited from the OS-specific detection mechanism, achieving superior detection performance compared to Nodlink, particularly on Linux, where system activities are highly dependent on command-line interactions. Unlike Nodlink, which employs a generic detection approach, our method adapts to system-specific behaviors by leveraging known attack signatures across different OS platforms. This tailored strategy enhances detection accuracy and reliability, ensuring more precise identification of malicious activities.
IV-F3 Adaptive Detection in Dynamic Cyber Environments: Minimizing False Positives
Figure 13(b) shows the node-level $F_{1}$ score before and after the use of the adaptive interface across different OS platforms. The results of the worst-case and best-case scenarios are derived from a series of fine-tuning attempts conducted during our evaluation.
It can be observed that, AI detectors are highly sensitive to varying operational environments, necessitating dynamic fine-tuning to adapt effectively to changing conditions. Similarly, each OS environment requires fine-tuning to achieve optimal detection results.
However, most research efforts [17, 7, 35, 5, 36] fail to provide a user-friendly, adaptive interface for fine-tuning their models. Instead, they rely on pre-defined parameters within their detection mechanisms, resulting in suboptimal performance in real-world detection scenarios. Our adaptive interface effectively addresses this challenge, ensuring that our solution can be seamlessly applied in real-world detection scenarios.
(a) GPT Recommender Results on E3-TRACE
(b) GPT Recommender Results on E3-THEIA
(c) GPT Recommender Results on E3-CADETS
(d) GPT Recommender Results on Real-Time Attack Scenario #1
(e) GPT Recommender Results on Real-Time Attack Scenario #2
(f) GPT Recommender Results on Real-Time Attack Scenario #3
Figure 14: Comparison of GPT recommender results across E3 datasets and real-time attack scenarios.
(a) Progress Injection and Command-Line Evasion
(b) Living off the Land Binary (LOLBin)
Figure 15: Visualization of (a) Process Injection and Command-Line Evasion (Silent Lynx [30, 31]); (b) Living off the Land Binary (LOLBin Evilnum [28]); with GPT recommender.
(a) Advanced Command Transformation
(b) Backdoor Execution
(c) Base64 Shell Command Obfuscation
(d) Log Deletion
(e) Timestamp Forgery
(f) Firefox Extension Attack
Figure 16: Visualization of the differences in detected command-line execution chains between DefendCLI and Nodlink [17], based on the E3-series datasets and real-time evaluation.
IV-F4 Enhancing the Practicability of our Solution for Actionable Attack Provenance Analysis
We replace the traditional alarm scoring framework with a GPT-powered recommender system to prioritize critical alarms, effectively reducing alarm fatigue and improving response efficiency. Figure 15 shows the visualization of some InfoPaths that are detected in our real-time evaluation with the GPT recommender. The GPT recommender first analyzes and explains the intent behind each command-line that appears potentially malicious, and then prioritizes the alarm for further examination.
Figure 14 presents the overall results of the GPT recommender across E3 datasets and real-time attack scenarios. The GPT recommender generates the top-5 malicious snapshot alarms for initial examination, while the remaining alarms are restored for further review. If the top-5 snapshot alarms exhibit malicious traces, security analysts then proceed to examine the other alarms to gather additional evidence.
IV-G The Improvement of Command-Line Detection Capability
Table IX presents a comparison of the scale of attack-related command-line activities detected by DefendCLI and Nodlink [17]. It can be observed that our solution demonstrates superior overall effectiveness in identifying a greater number of malicious command-line activities.
$\blacksquare$ How well does DefendCLI in the detection of obfuscation compared to Nodlink? Figures 16(a) and 16(c) illustrate the detection results of both detectors in handling advanced command-line transformations and Base64-encoded shell command obfuscation.
In Figure 16(a), the string 732F5B202F67732F5B...202F67 represents a transformation of an ASCII command into its hexadecimal form as a defense evasion technique. When converted back to ASCII, it reveals the command $s/[[:blank:]]\textbackslash{}+//g$, which is used to clean up injected spaces in obfuscated shellcode designed to evade detection.
In Figure 16(c), the string 2E2F676C20262E2F67...6C2026 represents a base64 obfuscated command-line for defense evasion. When decoded the content, it reveals the command $gtcache~\&\textgreater{}/dev/null~\&$, which runs a backdoor process in the background, suppressing output and ensuring persistent session activity without logging.
Based on our evaluation results, DefendCLI successfully detects advanced command-line attacks, whereas Nodlink [17] fail to identify them. This demonstrates the superior capability of DefendCLI in handling complex command-line obfuscation techniques, ensuring more comprehensive attack detection.
$\blacksquare$ How well does DefendCLI in the attack correlation with the low-frequency event compared to Nodlink? Figures 16(b), 16(d), 16(e), and 16(f) illustrate the detection results of both detectors in handling complex attack-chain correlations involving low-frequency events.
It can be observed that DefendCLI exhibits superior overall effectiveness in identifying and correlating a broader range of relevant command-line activities, improving the analysis of the provenance of the attack.
In contrast, Nodlink [17] fails to detect the command chmod+xtcexecchmod~+x~tcexec for escalating the backdoor privileges of tcexectcexec as shown in Figure 16(b). It also does not identify clear_consoleclear\_console as the root cause of the attack as shown in Figure 16(d) and is unable to reconstruct the full sequence of attack activities due to the absence of the timestamp forgery command $date-dnull+\%s$ as shown in Figure 16(e). While it detects the crashed execution of the Firefox browser, it fails to correlate subsequent actions involving the updatemodeupdate-mode script, which signifies a deeper attack footprint related to information modification as shown in Figure 16(f).
IV-H Runtime Performance
Table X presents the execution performance of DefendCLI across diverse datasets, highlighting its efficiency through parallelization. For the E3 data sets, DefendCLI runs on a 64,GB RAM, 16-core virtual machine, processing approximately 5,000 records in 6 seconds using multiprocessing. For industrial-scale, real-time detection tasks (A1, A2, A3), the system is deployed on Microsoft Azure DataLake [21], leveraging Kafka-based streaming and ETL pipelines. Under this setup, DefendCLI achieves end-to-end response times of roughly 2 seconds, including network latency, for high-volume threat detection workloads.
TABLE IX: The Scale of Attack-Related Command-Line Detected by DefendCLI and Nodlink
Dataset DefendCLI Nodlink Potential Weakness
E3-TRACE 18% 11% Obfuscation
E3-THEIA 19% 17% Context Awareness
E3-CADETS 18% 13% Low-Frequency Event
A1 15% 8% Context Awareness
A2 25% 12% Low-Frequency Event
A3 9% 3% Low-Frequency Event
TABLE X: Runtime Performance
Dataset Runtime Batch Size Platform
E3-TRACE $\approx$ 5 secs 5,000 records 64GB RAM - 16 Cores
E3-THEIA $\approx$ 5 secs 5,000 records 64GB RAM - 16 Cores
E3-CADETS $\approx$ 7 secs 5,000 records 64GB RAM - 16 Cores
A1 A2 A3 $\approx$ 2 secs / Microsoft Azure DataLake
V Conclusion
In conclusion, DefendCLI addresses critical gaps by introducing a multilevel command-line-driven differentiation framework that enhances attack provenance analysis. By leveraging graph algorithms, explainable AI, and multi-stage correlation, DefendCLI refines provenance analysis to detect attacks existing methods often miss. Unlike traditional models relying on pre-trained normal behavior profiles, it provides contextually aware anomaly detection, improving precision while reducing irrelevant alerts.
Through an extensive evaluation of real-world APT threats, DefendCLI demonstrates effectiveness in identifying sophisticated attack chains, detecting overlooked threats, and improving actionable intelligence for security teams. By enhancing reliability, reducing detection noise, and increasing context awareness, it delivers a practical, scalable, high-precision EDR solution to strengthen cybersecurity defenses.